OpenStudy (anonymous):

Which ratio is equal to 4/7?

OpenStudy (anonymous):

Options?

OpenStudy (jdoe0001):

one should note that \(\large \cfrac{4}{7}\) is in itself a ratio already

OpenStudy (anonymous):

14/17 41/71 8/11 8/14

OpenStudy (whpalmer4):

Okay, if you multiply 4/7 by the same number both top and bottom, is there any such number that will give you one of your answer choices?

OpenStudy (anonymous):

8/14 = 2.4/2.7 = 4/7

OpenStudy (whpalmer4):

That's rather confusing as written, @zairhenrique \[\frac{8}{14} = \frac{2*4}{2*7} = \cancel{\frac{2}{2}}*\frac{4}{7} = \frac{4}{7}\]

OpenStudy (whpalmer4):

(I assume you meant a center dot to signify multiplication, but here it looks like a decimal point!)